Fermi National Laboratory
 

Accelerator Update for July 10, 2006 - July 12, 2006

 

Monday July 10

The day shift began with Operations monitoring store 4818 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) in access, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, with NuMI off due to a horn cooling problem, and with MiniBooNE taking beam.

At 8:25 AM, a Tevatron (TeV) RF station (TRF8) tripped off on an over current indication. Operations contacted an expert.

Pbar came out of access at 9:48 AM.

Operations established beam to the Meson MTest experiment at 9:53 AM.

At 10:13 AM, Operations reported that an Accumulator quadrupole power supply (A:LQ) would not ramp. An expert investigated and discovered a problem with its contactor. He completed the repair about an hour later.

Pbar began stacking at 11:10 AM.

At 11:34 AM, Operations reported that A:LQ had its output current drop, which caused the loss of some antiprotons. An expert investigated.

The power supply expert reported at 1:30 PM that A:LQ was repaired.

Pbar resumed stacking at 1:52 PM.

An RF expert reported at 2:58 PM that TRF8 had suffered a power amplifier spark, but it should turn on okay.

At 6:43 PM, Operations reported that the Linac Klystron Debuncher had tripped on a cavity LCW flow for the first of five times in about forty-five minutes. Operators reset it each time and then after the last trip added a little more pressure to the expansion tank.

Operations began preparing the Recycler at 7:06 PM to stash antiprotons. The stash ended and Pbar resumed stacking about 8:24 PM. MTest resumed taking beam a few minutes later.

MTest experimenters reported that they were through with the beam for the night at 9:13 PM.

Recycler began preparing to stash around 11:45 PM.

Tuesday July 11

The midnight shift began with Operations monitoring store 4818 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

The stash ended at 12:10 AM.

At 2:12 AM, Operations reported that there was a problem with the Recycler’s Pelletron. Operations called an expert who came in and was able to get it on in time for the next stash.

The Recycler began preparing to stash at 4:25 AM. After some transfer problems, operators completed the stash. Pbar resumed stacking at 5:44 AM.

MTest began taking beam around 9 AM.

Operations terminated store 4818 at 9:09 AM.

The Cryo Coordinator reported at 9:28 AM that cryo system technicians were beginning work on the TeV sector F4 wet engine. They completed their work about an hour later.

At 9:49 AM, a RF expert turned off the TeV RF so that FESS techs could clean the sector F0 pond strainer. They completed their work within thirty minutes.

The Recycler began preparing to stash at 10:34 AM.

A power supply expert began maintenance on five Switchyard (SY) supplies at 10:34 AM. He completed his work about an hour and forty-five minutes later.

The stash ended and Pbar resumed stacking at 11:37 AM.

Operations put the TeV through a dry squeeze at 12:57 PM

Operations put the TeV into shot setup at 1:30 PM.

Operations established store 4821 at 4:08 PM. Its initial luminosity was 156.85E30.

The Recycler began preparing to stash at 6:31 PM. The stash ended and MTest resumed taking beam at 6:49 PM. Pbar resumed stacking at 6:55 PM.

MTest reported 9:04 PM that they were finished with the beam for the night.

The Recycler began preparing to stash at 11:24 PM. The stash ended near midnight and Pbar resumed stacking.

Wednesday July 12

The midnight shift began with Operations monitoring store 4821 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

At 12:03 AM, the TeV aborted store 4821 due to an indication of a beam position problem at CDF. There was no quench. Operations investigated, but could find no apparent problem. Experts thought the abort might have been caused by a BLM failure. An instrumentation expert came in and checked their equipment, but found nothing wrong.

The Recycler began preparing to stash at 4:16 AM. The stash ended and Pbar resumed stacking at 4:43 AM.

Operations put the TeV into shot setup at 6:35 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store. D0 has a detector problem. They will take time during the day to see if they can make detector repairs from outside of their hall. If D0 cannot make these repairs from outside their hall, tomorrow (7/13/06) at 7 AM they will begin a sixteen-hour access. If this access occurs, the Run Coordinator will allow the TeV, Pbar, Booster, and Linac to access their machines for two to three hours.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/12/2006    email Fermilab